OpenStudy (anonymous):

What is the x-intercept, the y-intercept, and the slope of y=3x+5

jimthompson5910 (jim_thompson5910):

The slope of y = mx+b is m y=3x+5 is in the form y = mx+b where m = 3 So the slope of y=3x+5 is 3

jimthompson5910 (jim_thompson5910):

To find the x intercept, plug in y = 0 and solve for x

jimthompson5910 (jim_thompson5910):

To find the y intercept, plug in x = 0 and solve for y

jimthompson5910 (jim_thompson5910):

Tell me what you get for each

OpenStudy (anonymous):

x= -5/3 ??

OpenStudy (anonymous):

y=5 ?

jimthompson5910 (jim_thompson5910):

you are correct on both, nice work

jimthompson5910 (jim_thompson5910):

so the x intercept is x = -5/3 the x-intercept as a point is (-5/3, 0)

OpenStudy (anonymous):

Thanks so much! have a great day! :)

jimthompson5910 (jim_thompson5910):

yw, you too